### v1

@@ 3.2 Learnability of Distorted Latent Representations
To explore this, we design the following experiment: diffusion models are known to overfit when trained or fine-tuned on a very small number of image samples. In such cases, the overfitted model exhibits memorization, generating samples nearly identical to the training images when given the same text prompts used during fine-tuning. We first fine-tune the pre-trained LDM on a single image obtained from one identity in the dataset. Once the model overfits, we generate an image from the fine-tuned model using the exact same text prompt. For the latent representation $z_{c}$ of a clean image, we expect the generated $\tilde{z}_{c}$ to be close to $z_{c}$ in the latent space. The difference ratio for this clean image is defined as $s_{c}=\|z_{c}-\tilde{z}_{c}\|/z_{b}$, where $z_{b}=\max(z_{c})-\min(z_{c})$ serves as a normalization term. Subsequently, the difference ratios for $s_{r}$, $s_{a}$ and $s_{a}^{(\text{cat})}$ are defined as $\|z_{r}-\tilde{z}_{c}\|/z_{b}$, $\|z_{a}-\tilde{z}_{c}\|/z_{b}$ and $\|z_{a}^{\text{cat}}-\tilde{z}_{c}\|/z_{b}$, respectively. For simplicity, we also define the $s_{r}$ range as $[s_{c}-\|s_{c}-s_{r}\|,s_{c}+\|s_{c}-s_{r}\|]$. We present the evaluation results in Figure 5. The difference ratios are evaluated for nine protection methods based on the sample clean image from an identity in the VGGFace2 dataset and two fine-tuning approaches, DreamBooth and LoRA. As shown in Figure 5, $s_{a}$ consistently falls within or near the range of $s_{c}\pm s_{r}$ across all protection methods and both fine-tuning approaches, which suggests that the latent representations of both clean samples and adversarial examples can be effectively learned by the diffusion process. This indicates that the primary reason for the effectiveness of adversarial examples as protective perturbation methods lies in the distortion of their latent representations, rather than the limitation of the diffusion process in learning these distorted representations.
In summary, this section provides a deeper investigation into why adversarial examples of LDMs are effective as protective perturbations against unauthorized data usage. We identify that the primary reason lies in the distortion of latent representations for these adversarial examples. Furthermore, we verify that these distorted latent representations can still be effectively learned by the diffusion process, confirming that the distortion itself is the key factor behind their effectiveness.
@@ 4.1 Contrastive Adversarial Loss
We have observed that adversarial examples are effective as protective perturbations because their latent representations are distorted. This indicates that the latent autoencoder $\{\mathcal{E}_{\phi},\mathcal{D}_{\omega}\}$ of the LDM is attacked. To mitigate this, we apply contrastive adversarial training to the latent autoencoder to realign the latent representations of adversarial examples. We first introduce the contrastive adversarial loss used for contrastive adversarial training as follows. Given the protected image $x_{a}$ and the latent autoencoder $\{\mathcal{E}_{\phi},\mathcal{D}_{\omega}\}$, we define the loss function as
$\mathcal{L}_{\text{cat}}(\mathcal{E}_{\phi},\mathcal{D}_{\omega},x_{a})=\mathbb{E}_{x_{a}}\|\mathcal{D}_{\omega}(\mathcal{E}_{\phi}(x_{a}))-x_{a}\|^{2}_{2}.$ (1)
@@ 4.2 CAT Adapters
Fine-tuning all layers of the latent autoencoder may affect the model’s performance in encoding and decoding unprotected image samples. Moreover, we aim to introduce additional parameters to expand the parameter space, improving the model’s resilience against adversarial examples after the contrastive adversarial training. Therefore, we incorporate adapters in our proposed contrastive adversarial training. Specifically, we apply the LoRA Hu et al. (2022) adapter to all convolutional and attention layers within the latent autoencoder. For the model weights $\phi_{0}\in\mathbb{R}^{d\times k}$ of the latent encoder, the latent representation $z^{(\text{cat})}_{a}$ of a protected sample $x_{a}$ is given by
@@
Once CAT is applied and the CAT adapters are obtained, customization on protected image samples is performed using the latent autoencoder augmented with the CAT adapters. For image generation, since the denoising module is fine-tuned on the realigned latent representation $z^{(\text{cat})}_{a}$, the generated output in the latent space for a given prompt is also realigned. Consequently, it is decoded into the image $\tilde{x}_{a}^{(\text{cat})}$ in the pixel space instead of $\tilde{x}_{a}$, effectively neutralizing the protective perturbations.
@@
5 Experiments
5.1 Experimental Setup
Evaluated Protective Perturbation Methods To evaluate the effectiveness of our proposed CAT method in compromising the robustness of protective perturbations, we test CAT across nine protective perturbation settings: (1) AdvDM(+) Liang et al. (2023), (2) AdvDM(-) Xue et al. (2023), (3) Mist Liang & Wu (2023), (4) SDS(+) Xue et al. (2023), (5) SDS(-) Xue et al. (2023), (6) SDST Xue et al. (2023), (7) Glaze Shan et al. (2023), (8) Anti-DreamBooth Van Le et al. (2023), and (9) Metacloak Liu et al. (2024). The (+) sign denotes the use of gradient ascent to construct the perturbation, whereas the (-) sign denotes the use of gradient descent, following the setting in Xue et al. (2023). We set the perturbation budget for all evaluated protection methods to $\delta=16/255$. Since Glaze does not release its codebase and its application does not allow adjustment of the specific perturbation budget, we use its maximum perturbation strength.
Datasets We evaluate the effectiveness of CAT in two main diffusion-based image synthesis applications: (1) object-driven image synthesis and (2) style mimicry. Object-driven image synthesis focuses on generating images of a specific object based on a given text prompt, such as the face of a particular person or a photo of a specific dog. Style mimicry, conversely, aims to generate images that mimic the style of the provided customization images.
For object-driven image synthesis, we focus on human face generation using the CelebA-HQ Karras et al. (2018) and VGGFace2 Cao et al. (2018) datasets. For each dataset, we select five identities, with 12 images per identity, equally split into three subsets: a clean reference subset, a target protection subset, and an additional clean reference subset. For style mimicry, due to page constraints, we present only qualitative results to demonstrate the effectiveness of CAT. We use a subset of Claude Monet’s artworks from the WikiArt dataset Mancini et al. (2018), consisting of 12 images, divided into three subsets following the same configuration. All images are center-cropped and resized to $512\times 512$.
@@ 5.1 Experimental Setup
We select all convolutional and attention layers within the module as the target layers. The adapter rank is set to $r=128$ for the CAT-both setting and $r=256$ for the CAT-en and CAT-de settings to maintain the same number of parameters. During CAT, only the weights of the attached adapters are updated, while all other weights remain frozen. Training is performed with a batch size of 4 and a learning rate of 1 ×$\times 10^{-4}$ for 1000 steps.
@@
Customization Settings For the customization settings, we adopt the widely used DreamBooth Ruiz et al. (2023) and LoRA Hu et al. (2022) methods, using the latest Stable Diffusion v2.1 as the pre-trained LDM. Details of the customization settings are provided in Appendix B.
@@
Evaluation Metrics For Qualitative Experiments The object-driven image synthesis task aims to generate human faces that preserve the identity features from the given protected images. Therefore, for qualitative experiments, we use Face Detection Rate (Retina-FDR), measured with the RetinaFace detector Deng et al. (2020), to evaluate whether a face is detected in the generated images. If a face is detected, we extract its embedding using the ArcFace recognizer Deng et al. (2019) and compute the cosine distance between the extracted embedding and those from the clean dataset, denoted as Identity Score Matching (ISM). To quantify the expected identity similarity in the generated image set, we define Face Similarity Score (FSS) as the product of Retina-FDR and ISM.
Moreover, we employ the TOPIQ method Chen et al. (2024), trained on the face IQA dataset and implemented in Chen & Mo (2022), to measure Face Detection Rate (TOPIQ-FDR) and Face Image Quality (FIQ). To quantify the expected quality of the generated face images, we define their product as the Face Quality Score (FQS). We use two text prompts for sampling, ”a photo of sks person” and ”a dslr portrait of sks person”. For each test prompt, we generate 30 images and compute the average metric scores across all tested identities to obtain the final results.
@@
5.2 Object-driven Synthesis Results
The evaluation results for applying CAT methods in the object-driven image synthesis task are presented in Table 1. We observe that both the CAT-both and CAT-en settings significantly improve task performance compared to the baseline, which directly customizes using the protected image samples across all evaluated protections. This demonstrates the effectiveness of these two settings in compromising the protective perturbations. Furthermore, the CAT-both setting achieves a higher FSS score than CAT-en, while CAT-en obtains a better FQS score overall. Additionally, the CAT-de setting has little to no effect on improving task performance, further verifying that the primary reason adversarial examples function effectively as protective perturbations is their latent representation distortion. Qualitative results for the identity n000050 from the VGGFace2 dataset, as shown in Figure 6 (for the identity 5 from the CelebA-HQ dataset in Section A.1), further confirm the observations from the quantitative results. Moreover, we also experimented with object-driven image synthesis customization using LoRA, as shown in Section A.2. The results verify the effectiveness of CAT in the LoRA customization setting across all nine evaluated protective perturbations and two datasets.
@@ 5.3 Style Mimicry Results
We evaluate only the CAT-both and CAT-en settings for style mimicry customization using DreamBooth to verify their effectiveness in compromising the protective perturbations. Qualitative results are presented in Figure 7, demonstrating that applying CAT leads to a significant improvement in the task performance of style mimicry customization. Across all evaluated protective perturbations, applying CAT preserves the specific artist’s style and improves the fidelity of generated images while reducing the distortions introduced by the protections. These results demonstrate that CAT weakens the robustness of existing protective perturbations, enabling more precise style mimicry and highlighting the need for more effective defense mechanisms.
5.4 Ablation Study: Impact of CAT Adapter Rank
To investigate the impact of the CAT adapter rank on the effectiveness, we conduct experiments using different rank settings on the VGGFace2 dataset under the AdvDM(+) protective perturbation and CAT-both setting. The results shown in Table 2 indicate that increasing the rank generally improves performance, with $r=128$ achieving the highest FSS (0.544) and FQS (0.460). This demonstrates that higher-rank adapters enhance robustness against protective perturbations.
However, higher-rank CAT adapters result in a larger model size. For instance, the CAT-both setting with $r=128$ produces a 104 MB adapter model, and this size continues to grow with increasing rank. This trade-off suggests that while higher-rank adapters enhance robustness against protective perturbations, practical deployment requires balancing performance gains and model efficiency.
Table 2: Effectiveness of the CAT-both method against AdvDM(+) protective perturbation on the VGGFace2 dataset at various CAT adapter rank settings for human face generation using DreamBooth. The optimal values for each metric are highlighted in bold.

### v2

@@ 3.3 Learnability of Distorted Latent Representations
To explore this, we design the following experiment: diffusion models are known to overfit when trained or fine-tuned on a very small number of image (Carlini et al., 2023). In such cases, the overfitted model exhibits memorization, generating samples that are nearly identical to the training image when given the same text prompt used during fine-tuning. We first fine-tune the pre-trained LDM on a single image obtained from one specific identity in the dataset. Once the model overfits, we generate an image from the fine-tuned model using the exact same text prompt. For the latent representation $z_{c}$ of a clean image, we expect the generated $\tilde{z}_{c}$ to be close to $z_{c}$ in the latent space. The difference ratio for this clean image is defined as $s_{c}=\|z_{c}-\tilde{z}_{c}\|/z_{b}$, where $z_{b}=\max(z_{c})-\min(z_{c})$ serves as a normalization term. Similarly, the difference ratios for $s_{r}$, $s_{a}$ and $s_{a}^{(\text{cat})}$ are defined as $\|z_{r}-\tilde{z}_{c}\|/z_{b}$, $\|z_{a}-\tilde{z}_{c}\|/z_{b}$ and $\|z_{a}^{\text{cat}}-\tilde{z}_{c}\|/z_{b}$, respectively. For simplicity, we also define the $s_{r}$ range as $[s_{c}-\|s_{c}-s_{r}\|,s_{c}+\|s_{c}-s_{r}\|]$. We present the evaluation results in Figure 5. The difference ratios are evaluated for nine protection methods based on the sample clean image from an identity in the VGGFace2 dataset and two fine-tuning approaches, DreamBooth and LoRA. As shown in Figure 5, $s_{a}$ consistently falls within or near the range of $s_{c}\pm s_{r}$ across all protection methods and both fine-tuning approaches, which suggests that the latent representations of both clean and adversarial samples can be effectively learned by the diffusion process. This indicates that the primary reason for the effectiveness of adversarial examples as protective perturbation methods lies in the distortion of their latent representations, rather than the limitation of the diffusion process in learning these distorted representations.
In summary, this section provides a deeper investigation of why adversarial examples of LDMs are effective as protective perturbations against unauthorized data usage. We identify that the primary reason lies in the distortion of latent representations for these adversarial examples. Furthermore, we verify that these distorted latent representations can still be effectively learned by the diffusion process, confirming that the distortion itself is the key factor behind their effectiveness. We emphasize that this conclusion is mainly based on empirical experimental observations, and other potential contributing factors may exist.
@@ 4.1 Contrastive Adversarial Loss
We have observed that adversarial examples are effective as protective perturbations because their latent representations are distorted. This indicates that the latent autoencoder $\{\mathcal{E}_{\phi},\mathcal{D}_{\omega}\}$ of the LDM is attacked. To mitigate this, we apply contrastive adversarial training to the latent autoencoder to realign the latent representations of adversarial examples. We first introduce the contrastive adversarial loss used for our adversarial training strategy as follows. Given the protected image $x_{a}$ and the latent autoencoder $\{\mathcal{E}_{\phi},\mathcal{D}_{\omega}\}$, we define the loss function as
$\mathcal{L}_{\text{cat}}(\mathcal{E}_{\phi},\mathcal{D}_{\omega},x_{a})=\mathbb{E}_{x_{a}\sim p(x_{a})}\|\mathcal{D}_{\omega}(\mathcal{E}_{\phi}(x_{a}))-x_{a}\|^{2}_{2}.$ (1)
@@ 4.2 CAT Adapters
Fine-tuning all layers of the latent autoencoder may affect the model’s performance in encoding and decoding unprotected image samples. Moreover, we aim to introduce additional parameters to expand the parameter space, improving the model’s resilience against adversarial examples after the contrastive adversarial training. Therefore, we incorporate adapters in our proposed contrastive adversarial training. Specifically, we apply the LoRA (Hu et al., 2022) adapter to all convolutional and attention layers within the latent autoencoder. For the model weights $\phi_{0}\in\mathbb{R}^{d\times k}$ of the latent encoder, the latent representation $z^{(\text{cat})}_{a}$ of a protected sample $x_{a}$ is given by
@@
Once CAT is applied and the CAT adapters are obtained, customization on protected image samples is performed using the latent autoencoder augmented with the CAT adapters. For image generation, since the denoising module is fine-tuned on the realigned latent representation $z^{(\text{cat})}_{a}$, the generated output in the latent space for the given text prompt is also realigned. Consequently, it is decoded into the image $\tilde{x}_{a}^{(\text{cat})}$ in the pixel space instead of $\tilde{x}_{a}$, effectively neutralizing the protective perturbations.
5 Experiments
5.1 Experimental Setup
Evaluated Protective Perturbation Methods To evaluate the effectiveness of our proposed CAT method in compromising the robustness of protective perturbations, we test CAT across nine protective perturbation settings: (1) AdvDM(+) (Liang et al., 2023), (2) AdvDM(-) (Xue et al., 2024), (3) Mist (Liang & Wu, 2023), (4) SDS(+) (Xue et al., 2024), (5) SDS(-) (Xue et al., 2024), (6) SDST (Xue et al., 2024), (7) Glaze (Shan et al., 2023), (8) Anti-DreamBooth (Le et al., 2023), and (9) Metacloak (Liu et al., 2024). The (+) sign denotes the use of gradient ascent to construct the perturbation, whereas the (-) sign denotes the use of gradient descent, following the setting in (Xue et al., 2024). We set the perturbation budget for all evaluated protection methods to $\delta=16/255$. Since Glaze does not release its codebase and its application does not allow adjustment of the specific perturbation budget, we use its maximum perturbation strength.
@@ 5.1 Experimental Setup
Datasets We evaluate the effectiveness of CAT in two main diffusion-based image synthesis applications: (1) object-driven image synthesis and (2) style mimicry. For object-driven image synthesis, we focus on human face generation using the CelebA-HQ (Karras et al., 2018) and VGGFace2 (Cao et al., 2018) datasets. For each dataset, we select five identities, with 12 images per identity, equally split into three subsets: a clean reference subset, a target protection subset, and an additional clean reference subset. For style mimicry, we use a subset of Claude Monet’s artworks from the WikiArt dataset (Mancini et al., 2018), consisting of 12 images, divided into three subsets following the same configuration. All images are center-cropped and resized to $512\times 512$.
@@
We select all convolutional and attention layers within the module as the target layers. The adapter rank is set to $r=128$ for the CAT-both setting and $r=256$ for the CAT-en and CAT-de settings to maintain the same number of added parameters. During CAT, only the parameters of the attached adapters are updated, while all other weights remain frozen. Training is conducted with a batch size of 4 and a learning rate of 1 ×$\times 10^{-4}$ for 1000 steps.
@@
Customization Settings For the customization settings, we adopt the widely used DreamBooth (Ruiz et al., 2023) and LoRA (Hu et al., 2022) methods, using the latest Stable Diffusion v2.1 as the pre-trained LDM. Details of the customization settings are provided in Appendix B.
Table 2: Quantitative results for object-driven image synthesis using CAT methods customized in DreamBooth, compared to Noisy-Upscaling (NU) and Gaussian Filtering (GF) methods. The table presents performance metrics FSS and FQS for the CelebA-HQ and VGGFace2 datasets across different protective perturbation settings. The optimal values among different CAT settings for each metric are bolded.
Dataset CelebA-HQ VGGFace2
Metric FSS $\uparrow$ FQS $\uparrow$ FSS $\uparrow$ FQS $\uparrow$
Method CAT-both CAT-en NU GF CAT-both CAT-en NU GF CAT-both CAT-en NU GF CAT-both CAT-en NU GF
AdvDM(+) 0.643 0.529 0.531 0.492 0.431 0.448 0.481 0.352 0.534 0.560 0.518 0.506 0.481 0.578 0.506 0.363
AdvDM(-) 0.623 0.571 0.469 0.607 0.549 0.611 0.498 0.526 0.564 0.547 0.529 0.563 0.635 0.676 0.563 0.506
Mist 0.572 0.501 0.491 0.488 0.597 0.580 0.475 0.493 0.557 0.521 0.566 0.518 0.662 0.701 0.518 0.437
SDS(+) 0.602 0.499 0.599 0.409 0.413 0.423 0.503 0.302 0.486 0.508 0.498 0.402 0.438 0.569 0.402 0.281
SDS(-) 0.678 0.599 0.468 0.583 0.597 0.587 0.494 0.493 0.570 0.569 0.509 0.593 0.700 0.671 0.593 0.558
SDST 0.594 0.485 0.470 0.446 0.587 0.588 0.474 0.464 0.559 0.546 0.521 0.538 0.627 0.671 0.538 0.482
Glaze 0.610 0.577 0.533 0.547 0.618 0.676 0.496 0.533 0.607 0.576 0.503 0.549 0.733 0.723 0.549 0.562
Anti-DB 0.662 0.597 0.540 0.575 0.608 0.664 0.469 0.543 0.584 0.546 0.566 0.548 0.636 0.656 0.548 0.499
MetaCloak 0.642 0.578 0.521 0.540 0.460 0.475 0.395 0.324 0.560 0.631 0.566 0.542 0.504 0.633 0.542 0.349
@@
Evaluation Metrics For Quantitative Experiments The object-driven image synthesis task aims to generate human faces that preserve the identity features from the given protected images. Therefore, for qualitative experiments, we use Face Detection Rate (Retina-FDR), measured with the RetinaFace detector (Deng et al., 2020), to evaluate whether a face is detected in the generated images. If a face is detected, we extract its embedding using the ArcFace recognizer (Deng et al., 2019) and compute the cosine distance between the extracted embedding and those from the clean dataset, denoted as Identity Score Matching (ISM) (Le et al., 2023). To quantify the expected identity similarity in the generated image set, we define Face Similarity Score (FSS) as the product of Retina-FDR and ISM. Moreover, we employ the TOPIQ method (Chen et al., 2024), trained on the face IQA dataset and implemented in (Chen & Mo, 2022), to measure Face Detection Rate (TOPIQ-FDR) and Face Image Quality (FIQ). To quantify the expected quality of the generated face images, we define their product as the Face Quality Score (FQS). We use two text prompts for sampling, ”a photo of sks person” and ”a dslr portrait of sks person”. For each test prompt, we generate 30 images and compute the average metric scores across all tested identities to obtain the final results. We also present evaluation results using FID (Heusel et al., 2017) and CLIP-IQA (Wang et al., 2023), which are two widely used metrics for assessing image quality.
5.2 Object-driven Synthesis Results
The evaluation results for applying CAT methods in the object-driven image synthesis task are presented in Table 1. We observe that both the CAT-both and CAT-en settings significantly improve task performance compared to the baseline, which directly customizes using the protected image samples across all evaluated protections. This demonstrates the effectiveness of these two settings in compromising the protective perturbations. Furthermore, the CAT-both setting achieves a higher FSS score than CAT-en, while CAT-en obtains a better FQS score overall. Additionally, the CAT-de setting has little to no effect on improving task performance. The potential reason behind this is that CAT-de only adds adapters to the decoder, which has a limited impact on realigning latent representations. Fine-tuning only the decoder is more challenging, as the diffusion model learns from distorted latents that are highly diverse, making accurate reconstruction difficult. The weaker performance of CAT-de compared to CAT-en and CAT-both further supports our observation that latent distortion is the key factor behind the effectiveness of adversarial noise as a protective perturbation. Qualitative results for one identity from the VGGFace2 dataset, as shown in Figure 6 (and for one identity from the CelebA-HQ dataset as shown in Figure 9 which is provided in Section C.3), further confirm the observations from the quantitative results.
Moreover, we conduct object-driven image synthesis customization experiments using LoRA as shown in Table 5, which is provided in Section C.1. The results verify the effectiveness of CAT in the LoRA customization setting across all nine evaluated protective perturbations and two datasets. The quantitative results measured by FID for customization using DreamBooth are also presented in Table 6, which is provided in Section C.2. These results further support the effectiveness of the CAT settings in compromising the existing protective perturbation.
@@ 5.2 Object-driven Synthesis Results
Table 3: Quantitative results for style mimicry using CAT methods customized in DreamBooth on the WikiArt dataset. The table reports CLIP-IQA scores, where higher values indicate better alignment with the target artistic style. The optimal values among different CAT settings for each metric are bolded.
CLIP-IQA $\uparrow$ Baseline CAT-both CAT-en
AdvDM(+) 0.343 0.390 0.621
AdvDM(-) 0.463 0.536 0.697
Mist 0.345 0.465 0.694
SDS(+) 0.285 0.366 0.366
SDS(-) 0.501 0.481 0.723
SDST 0.406 0.485 0.712
Glaze 0.532 0.614 0.730
Anti-DB 0.315 0.544 0.672
5.3 Comparison with Purification-Based Attacks
We compare our proposed CAT with two purification-based adaptive attacks against protective perturbations: Noisy-Upscaling (Hönig et al., 2025), which performs optimization-based purification, and traditional Gaussian filtering, which applies a low-pass filter to the perturbed images. Evaluations are conducted on both the CelebA-HQ and VGGFace2 datasets using CAT-both and CAT-en under the same experimental setting. For Noisy-Upscaling, we adopt the default configurations provided in the original paper. While for Gaussian filtering, we apply a Gaussian kernel of size 5 with $\sigma=1.0$. The quantitative results are presented in Table 2. We observe that CAT consistently achieves comparable or superior performance to both Noisy-Upscaling and Gaussian Filtering across all protective perturbations, in terms of both FQS and FSS. These results highlight the competitive effectiveness of CAT compared to existing purification-based methods.
@@ 5.4 Style Mimicry Results
We evaluate the CAT-both and CAT-en settings for style mimicry customization using DreamBooth. Quantitative results, measured by CLIP-IQA, are presented in Table 3. CAT consistently outperforms the baseline across all evaluated protection methods, achieving higher CLIP-IQA scores and confirming the effectiveness of our approach. Qualitative comparisons are shown in Figure 7, illustrating that CAT significantly improves task performance in style mimicry customization. Across all evaluated protection perturbations, applying CAT preserves the stylistic characteristics of the target artist while improving image fidelity and mitigating the distortions introduced by protections. These results demonstrate that CAT weakens the robustness of existing protection perturbations in style mimicry, highlighting the need for more advanced protection methods.
Table 4: Effectiveness of the CAT-both method against AdvDM(+) protective perturbation on the VGGFace2 dataset under various CAT adapter rank settings for human face generation using DreamBooth. The optimal values for each metric are bolded.
@@
5.5 Ablation Study: Impact of CAT Adapter Rank
To investigate the impact of the CAT adapter rank on the effectiveness, we conduct experiments using varying rank settings on the VGGFace2 dataset under the AdvDM(+) protective perturbation and CAT-both setting. The results shown in Table 4 indicate that increasing the rank generally leads to better performance, with $r=128$ achieving the highest FSS (0.544) and FQS (0.460). This demonstrates that higher-rank adapters enhance robustness against protective perturbations. However, higher-rank CAT adapters result in a larger model size. For instance, the CAT-both setting with $r=128$ yields a 104 MB adapter model, and this size continues to increase with higher ranks. This trade-off suggests that while higher-rank adapters enhance robustness against protective perturbations, practical deployment requires balancing performance gains and model efficiency.